The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                         http://www.astro.uiuc.edu/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 24 August 31, 1998

Contents

Conference summary and notes  1
Abstracts of 11 refereed papers  6
Abstract of 4 non-refereed papers 13

News

IAU Symposium 190 “New Views of the Magellanic Clouds”

The IAU Symposium 190 “New Views of the Magellanic Clouds” was held at Victoria, BC, Canada on July 13–17. The symposium started with Paul Hodge’s overview and ended with Sidney van den Bergh’s closing summary. During these five days many exciting scientific results were reported. To share these with readers who were not able to attend the meeting, we have prepared a summary. As we have available only sketchy notes and imperfect memory, this summary is inevitably subjective and may have inadvertently left out many important results reported in the meeting. We apologize to everyone whose results may not be adequately covered, or presented incorrectly. We refer the readers to the symposium proceedings for a complete and accurate coverage of the meeting.

Interstellar Medium. New telescopes and instruments made it possible to study the multiple phases of the ISM. High-dispersion UV spectroscopic observations made with IUE, HST GHRS/STIS, and ORFEUS-II in particular provide a great wealth of information on molecular, atomic, and ionized interstellar gases at different temperatures (de Boer, Marx-Zimmer, Richter). Several surveys of the interstellar gas in the MCs have been completed or are underway. The UM/CTIO Magellanic Cloud Emission-Line Survey takes images in Hα, [O III], [S II] lines and green and red continua. This survey will provide not only flux-calibrated catalogs of HII regions and planetary nebulae but also sensitive surveys of supernova remnants and emission-line stars (C.R. Smith). ROSAT X-ray observations have been used to analyze the hot, ionized medium in the Magellanic Clouds; the plasma temperature in the LMC varies from ∼ 10^{6.6} on the west to ∼ 10^{6.9} on the east; the plasma density is N_{e} ∼ 0.002 cm^{−3}. The SMC is not as bright as the LMC in X-rays; large-scale diffuse emission in the SMC is marginally detected only in the 1/4 keV band (Snowden). The ATCA HI survey of the MCs has been completed. A large number of HI shells and holes are detected. The HI of the LMC has a relatively uniform distribution with no concentration along the stellar bar. The HI disk has a scale height of ∼ 300 pc in the central region and flares out to 400 pc at 4 kpc from the center (Staveley-Smith, Kim, Stanimirovic). Analyses of HI 21-cm emission and absorption lines show that the cool HI in the Clouds is cooler than that in the solar neighborhood, which can be explained by the lower metallicity and dust-to-gas ratios in the Clouds (Dickey). A patchy hot 10^{5} K gas halo of the LMC has been demonstrated by the HST GHRS observations of interstellar CIV absorption lines (Wakker). A high resolution CO survey of the LMC has been completed with the NANTEN 4m telescope. Comparisons between the distributions of clusters and CO clouds indicate that young clusters are formed in CO clouds, but the CO clouds dissipate within a timescale of < 10^{7} yr (Fukui, Yamaguchi). The CO in the MCs can survive photodissociation only if they are shielded in dense clouds. The conversation factor N(H2)/I_{CO} varies strongly with physical conditions of individual clouds. The LMC conversion factor is a factor of 3 higher than the Galactic value (Rubio, Pak). The isotopic ratios among CO (^{13}C^{16}O, 13C^{18}O,^{12}C^{16}O,^{12}C^{18}O) provide critical tests for models of stellar evolution and star formation history (Chin, Heikkilä).

The MCs are excellent sites for studying interactions between stars and the interstellar medium. Multiple-wavelength observations of LMC supernova remnants (SNRs) have been analyzed to study their energetics, evolution, and interstellar environmental effects. X-ray observations by ROSAT and ASCA are useful in analyzing the plasma temperatures and abundances, and even the supernova types (Petre, Williams, Hwang, Dickel, Chu, Saito). The youngest SNR is SN1987A, in which supernova ejecta are beginning to impact the rings (Sonneborn). Superbubble structure, evolution, and statistics were comprehensively reviewed by Oey. Palous suggests that the interactions between massive stars and the interstellar medium and the interactions between shells produce the bubble-dominated interstellar medium in the LMC, as seen in the HI maps (Kim, Staveley-Smith). Supergiant shells of kpc sizes have been detected in X-rays, indicating the existence of hot plasma; the supergiant shells LMC2 and LMC4 have been studied in detail (Bomans, Caulet, Points). The 3-D and velocity structure of the interstellar medium near SN1987A has been analyzed, using the light echo and echelle observations (Crotts)

IMFs. Massey summarized evidence that despite a metallicity difference of a factor of 10 the upper IMFs of coeval populations (OB associations) in both Clouds and in the Milky Way agree within the errors (slopes of −1.3 to −1.4). The IMF of young MC clusters is in good agreement with a Salpeter slope (S. Beaulieu, Elson). Evidence for massive star formation in the field comes from, e.g., the presence of O3 stars (≈ 1 Myr) which must have formed there (Massey).

In the 30 Dor starburst region a normal Salpeter IMF is found for stars > 2.8 M_{⊙}. The numerous most massive stars (O3) appear to have formed last (Hunter). Spectacular NICMOS images show the continued formation of massive, still embedded single and multiple O stars in the vicinity of R 136 (Walborn, Barbá). The IMF of R 136 below 3 M_{⊙} becomes flatter (Sirianni) and is dominated by pre-main sequence stars of ages from 1 to 10 Myr. A major NICMOS effort aimed at the low-mass IMF is detecting pre-main-sequence stars down to 0.2 M_{⊙} in 30 Dor region (Zinnecker, Brandner).

Models. New evolutionary tracks for rotating stars and binary stars show wider and brighter main sequences with 20%–30% longer main-sequence life times. The new models reproduce the blueto-red supergiant ratio as a function of metallicity and account for the observed nitrogen abundance variations (Langer).

Abundances. Garnett reviewed H II region abundances. He showed that 30 Dor is very homogeneous in abundance, indicating that temperature fluctuations and supernovae have minor effects on abundance determinations. Nebular compositions and stellar abundances agree very well for nonrefractory elements except C,N. A metallicity gradient in the LMC is still an open question.

V. Smith reviewed the chemical enrichment history of the MCs as evidenced by α and Fe-peak elements as well as s- and r-process species.

The mean abundance of the young SMC population is [Fe/H] = −0.7 dex, and −0.3 dex for the LMC. Metallicity differences between young clusters and young field stars are small. The mean stellar oxygen and α abundances are [O/Fe] = −0.18_{SMC}, −0.15_{LMC}, −0.3_{MW}, and [α/Fe] = −0.04_{SMC}, −0.15_{LMC} (Hill). O abundances in the ISM and B V, B I to K I stars as well as Fe abundances in A I to M I supergiants show remarkable agreement. Both Clouds show little enrichment in C,O but a large range in N abundances (Venn).

Field populations. Westerlund reviewed the field star populations of the Magellanic Clouds and described a scenario of interaction-triggered bursts of star formation. The on-going large-scale photometric and microlensing surveys will lead to a comprehensive picture of the star formation history (Zaritsky, Harris).

SMC: Ages increase with increasing distance from the center of the SMC. The velocity dispersions of different populations are all very similar, and there is no clear rotational signature. The outer wing shows significantly higher velocity dispersions. The majority of older stars have ages of 6.3 to 9.5 Gyr, while no strong intermediate-age component of 2 – 4 Gyr (as in the LMC) is observed. A significant episode of star formation occurred 8 or 9.5 Gyr ago. The current star formation rate (SFR) is at least a factor of 2 lower. The mean SFR over a Hubble time is 0.09 M_{⊙} yr^{−1} in the SMC and 0.4 – 0.6 M_{⊙} yr^{−1} in the LMC (Hatzidimitriou).

LMC: HST color-magnitude diagrams of field star populations across the disk are quite similar and show a non-episodic, roughly constant SFR with an increase by a factor of 3 ≈ 2 Gyr ago as opposed to the age gap seen for intermediate-age clusters. Deep HST data further support a normal IMF in the LMC disk (Gallagher, Holtzman, Cole). The large-scale field star formation history shows evidence for long-lived stationary (≈ 200 Myr) chains of star formation in the LMC disk and migratory patterns along the bar (Grebel).

Clusters. Old globular clusters: Olsen and Johnson demonstrated that there is no distinguishable age difference between the oldest Magellanic and Milky Way halo globulars. The age range is no greater than 1 Gyr. Magellanic globulars are very similar to the outer Galactic halo clusters (Olszewski), although the outermost Galactic globulars beyond the Magellanic Clouds may be up to 2 Gyr younger (Hesser). The oldest LMC globulars appear to be in a disk-like system (Olszewski, Da Costa). The census of old (> 10 Gyr) clusters is fairly complete (Geisler).

Intermediate-age clusters: Both Magellanic Clouds show an “abundance gap” between their oldest and intermediate-age clusters. After a rapid initial enrichment (≈ 3 − 4 Gyr) metallicities increased more slowly. Intermediate-age clusters show a wide range of abundances (Da Costa). A good fit to the observed age-abundance data can be obtained with the bursting star formation model by Pagel & Tautvaišienė (Mighell). The discovery of three LMC clusters in the intermediate-age gap reduces the discrepancy between cluster and field age distribution (Sarajedini).

Young clusters: The age distribution of young (< 1 Gyr) clusters in both Clouds peaks at 100 – 200 Myr, coincident with predicted close encounters between the Clouds and the Milky Way. Recent field star formation and recent cluster formation trace each other very well. Large surveys triple the number of clusters, contributing mainly to the faint end of the cluster luminosity function (Grebel). The range of core radii increases with age. Tidal truncation does not occur before 1 Gyr (Elson). Young clusters may show evidence for dynamical mass segregation (Fischer). About half of the binary clusters in the Clouds may be physically connected pairs (Dieball).

Associations: OB associations show hierarchical structure, propagating star formation, and are 3× more numerous than previously assumed. Typical sizes are ≈ 250 pc and ≈ 600 pc (Kontizas, Maragoudaki). Propagating star formation is also seen in supergiant shell LMC 4 (Efremov).

Stars. IR surveys strongly increase the census of AGB, M, and C stars (Loup, Cioni, van Loon). 11,000 C stars are now known in the LMC, and 3,600 in the SMC. The number of C stars increases with metallicity. Fainter C stars are found in more metal-poor galaxies (Azzopardi).

SMC PNe are on average more compact than LMC PNe, possibly due to metallicity-related massloss. He-burning PNe (20% at end of AGB evolution) tend to be more extended (slower evolution) than H-burning PNe (80%) (Dopita).

The microlensing surveys (EROS, MACHO, OGLE) are completing the variable-star census in the Magellanic Clouds and lead to well-sampled lightcurves for a huge number of stars (Alves, J.-P. Beaulieu, Welch). There are more short-period Cepheids and more s-mode pulsators in the SMC (Marquette) than in the LMC. The metallicity dependence of P-L and P-R relations for Cepheids was discussed by Bono and Marconi.

Distances Highly accurate, direct distance measurements can be obtained with infrared surface brightness fluctuation methods, which are insensitive to metallicity and reddening. These methods allow to study also the tilt of the disk and depth extent. The resulting distance modulus for the LMC is 18.5 (Gieren). The Hipparcos Cepheid distance remains ≈ 0.2 mag larger, while the RR Lyr distance is ≈ 0.2 mag closer (Feast). Age and metallicity dependence of the red clump (Girardi, Cole) need to be thoroughly explored to make it a reliable distance indicator (Feast). The LMC distance based on a new analysis of SN 1987A is ≈ 18.6 (Panagia).

Dynamics/tidal interactions: N-body simulations predict the last close encounter between LMC and SMC ≈ 200 Myr ago, which produced the bridge/tail structure. The previous close encounter plus perigalacticon was 1.5 Gyr ago. The star formation rate in the SMC is enhanced by each tidal interaction. The N-body simulations also nicely reproduce the observed large-scale star formation patterns in the LMC affected by the off-centered bar (Gardiner). An unsolved puzzle are the abundances (−1.1 dex, Rolleston) seen in B V stars in the Magellanic bridge, which are much lower than expected from a recent SMC origin. A modified TreeSPH code has been used to simulate the MW-LMC-SMC interaction, and the calculations can easily produce the tidal features, high velocity clouds, the elongated geometry of the SMC, and the spiral pattern in HI gas and the stellar bar in the LMC (Li).

Spectacular data were presented by Putman showing Hi tidal tails between the Clouds and the Milky Way including leading arm features. Kinematic data for carbon stars trace SMC stars 4^{∘} to 20^{∘} from the LMC. A ring of material originating from the SMC is trapped around the LMC as massive perturber (Kunkel, Demers). A second red clump in outlying LMC populations indicates an SMC tidal feature in superposition (Geisler). A large-area survey for red giants around the Magellanic Clouds shows stars at distances expected for tidal debris from the Clouds (Majewski). Radial velocity dispersions for stars along the vertical extension of the red clump indicate that stars of Magellanic Cloud origin may now be located between us and the Clouds (Zaritsky), but the extension can also be explained by normal evolutionary effects (J.-P. Beaulieu). Both effects may be at work.

Eva Grebel & You-Hua Chu

What is 29 Doradus?

John R. Dickel^{1}

1 Astronomy Department, University of Illinois

At IAU Colloquium 190 on the Magellanic Clouds held in Victoria BC, 12-17 July 1998, I challenged the assembled multitude with the question

“What is 29 Doradus?”

You give me:

1) kind of object it is

2) coordinates (preferably J2000)

3) a reference so I can check it I give you: A bottle of Dickel Tennessee Sipping Whiskey

After several false leads, I am pleased to report on 29 July 1998 - Congratulations to Jim Kaler assisted by Al Calder who took me over to the UI Rare Book Room

and showed me page 90 of the publication “General Description and Information on the Stars” by Johann Bode 1801.

29 Doradus RA=84^{∘}15′33′′ Dec=−66^{∘}40′04′′ (1801) and m=6 Precession to J2000 gives RA=05^{h}37^{m}02^{s} Dec=−66^{∘}33′24′′ = HR1960 = HD37935 which is a

B9.5Ve star with a visual magnitude of 6.3

The identification allows the more accurate coordinates of RA=05^{h}39^{m}59.^{s}8 and Dec=−66^{∘}33′37′′ and the Hipparcos parallax gives a distance of 255 pc.

Second prize goes to Ed Olszewski and Ron Webbink who steered us to the correct publication. Another second prize goes to Michel Dennefeld, the first IAU Symposium 190 attendee to find the

correct answer, just one day after Kaler and Calder. He used the copy of Bode’s catalog at the Paris Observatory.

Third prize goes to Doug Welch who steered Ed Olszewski in the right direction. Some people have suggested that there are earlier catalogs which Bode may have copied but no

conclusive evidence has emerged yet. Bode lists 30 Dor as nebular and it precesses to about 4.′5 southwest of R136 which is about the center of the whole optical complex. I note that Bode also published an atlas which labels the Large and Small Nebulae with no mention of Magellan.

Many thanks to all and I hope others had as much fun in the search as I did.

Abstracts of Refereed Papers

Tidal disruption of the Magellanic Clouds by the Milky Way

M.E. Putman^{1}, B.K. Gibson^{1}, L. Staveley-Smith^{2}, G. Banks^{3}, D.G. Barnes^{4}, R. Bhatal^{5}, M.J. Disney^{3}, R.D. Ekers^{2}, K.C. Freeman^{1}, R.F. Haynes^{2}, P. Henning^{6},

H. Jerjen^{1}, V. Kilborn^{4}, B. Koribalski^{2}, P. Knezek^{7}, D.F. Malin^{8}, J.R. Mould^{1}, T. Oosterloo^{9}, R.M. Price^{2}, S.D. Ryder^{10}, E.M. Sadler^{11}, I. Stewart^{2}, F. Stootman^{5},

R.A. Vaile^{5†}, R.L. Webster^{4}, A.E. Wright^{2}

1Mount Stromlo & Siding Spring Obs., Australian National University, Weston Creek P.O., Weston, ACT 2611, Australia 2Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 3University of Wales, Cardiff, Department of Physics & Astronomy, P.O. Box 913, Cardiff CF2 3YB, Wales, U.K. 4University of Melbourne, School of Physics, Parkville, Victoria 3052, Australia 5University of Western Sydney Macarthur, Department of Physics, P.O. Box 555, Campbelltown, NSW 2560, Australia 6University of New Mexico, Department of Physics & Astronomy, 800 Yale Blvd. NE, Albuquerque, NM 87131, USA 7The Johns Hopkins University, Dept. of Physics & Astronomy, 34^{th} & N. Charles Streets, Baltimore, MD 21218, USA 8Anglo-Australian Observatory, P.O. Box 296, Epping, NSW 2121, Australia 9Istituto di Fisica Cosmica, via Bassini 15, I-20133, Milano, Italy 10Joint Astronomy Center, 660 North Aohoku Place, Hilo, HI 96720, USA 11University of Sydney, Astrophysics Department, School of Physics, A28, Sydney, NSW 2006, Australia †deceased

Interactions between galaxies are common and are an important factor in determining their physical properties such as position along the Hubble sequence and star-formation rate. There are many possible galaxy interaction mechanisms, including merging, ram-pressure stripping, gas compression, gravitational interaction and cluster tides. The relative importance of these mechanisms is often not clear, as their strength depends on poorly known parameters such as the density, extent and nature of the massive dark halos that surround galaxies. A nearby example of a galaxy interaction where the mechanism is controversial is that between our own Galaxy and two of its neighbours – the Large and Small Magellanic Clouds. Here we present the first results of a new Hi survey which provides a spectacular view of this interaction. In addition to the previously known Magellanic Stream, which trails 100^{∘} behind the Clouds, the new data reveal a counter-stream which lies in the opposite direction and leads the motion of the Clouds. This result supports the gravitational model in which leading and trailing streams are tidally torn from the body of the Magellanic Clouds.

Accepted by: Nature

For preprints, contact putman@mso.anu.edu.au
Also available from the URL http://msowww.anu.edu.au/~putman/pubs.html

The Large-Scale HI Structure of the Small Magellanic Cloud

S. Stanimirovic^{1,2}, L. Staveley-Smith^{2}, J.M. Dickey^{3}, R.J. Sault^{2}, and S.L. Snowden^{4}

1University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 2Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 3University of Minnesota, 116 Church St. SE, Minneapolis, MN 55455, USA 4Code 662, NASA/Goddard Space Flight Center, Greenbelt, MD 20771, USA

We combine new Parkes telescope observations of neutral hydrogen (HI) in the Small Magellanic Cloud (SMC) with an Australia Telescope Compact Array (ATCA) aperture synthesis mosaic to obtain a set of images sensitive to all angular (spatial) scales between 98 arcsec (30 pc) and 4^{∘} (4 kpc). The new data are used to study the HI spatial power spectrum over a range of contiguous scale sizes wider than those previously achieved in any other galaxy, including our own. The spatial power spectrum closely obeys the relation, P (k) ∝ k^{γ}, with γ = −3.04 ± 0.02, similar to values obtained by other authors for our own Galaxy which are in the range γ = −3.0 to −2.8. This is surprising given the very different morphology, gas-richness, star-formation rate and evolution of the two systems, and may imply similar mechanisms for structure formation. One interpretation of the P (k) power-law is that the interstellar medium (ISM) of the SMC is fractal in nature, consisting of a hierarchy of HI cloud structures created, for example, by homogeneous turbulence. The projected fractal dimension of D_{p} = 1.5 is similar to values obtained by other authors for molecular clouds in the Galaxy in the size range ∼ 0.05 to 100 pc. Such a model is consistent with a low space-filling factor for the neutral gas.

A kinematic study of the HI data reveals the existence of three supergiant shells which were previously undetectable in the ATCA data alone. These shells have diameters up to 1.8 kpc and require energies (in the standard supernova-driven models) up to 2 × 10^{54} erg. The structure and evolution of the ISM in the SMC is heavily influenced by the formation of these supergiant shells.

Accepted by: Monthly Notices of The Royal Astronomical Society

For preprints, contact sstanimi@atnf.csiro.au
Also available from the URL http://www.atnf.csiro.au/research/smc h1/

ORFEUS II echelle spectra: Absorption by H_{2} in the LMC^{∗}

K.S de Boer^{1}, P. Richter^{1}, D.J. Bomans^{2}, A. Heithausen^{3}, J. Koornneef^{4}

1 Sternwarte, Universität Bonn, Germany 2 Astronomisches Institut, Ruhr-Universität Bochum, Germany 3 Radioastronomisches Institut, Universität Bonn, Germany 4 Ruimteonderzoek, Kapteyn Institute, Groningen, The Netherlands

We present the first detection of H_{2} UV absorption profiles on the line of sight to the LMC. The star LH 10:3120 in the LMC was measured with the ORFEUS telescope and the Tü/HD echelle spectrograph during the space shuttle mission of Nov./Dec. 1996.

16 absorption lines from the Lyman band are used to derive the column densities of H_{2} for the lowest 5 rotational states in the LMC gas. For these states we find a total column density of N(H_{2})_{total} = 6.6 × 10^{18} cm^{−2} on this individual line of sight.

We obtain equivalent excitation temperatures of T_{ex} ≤ 50 K for the rotational ground state and T_{ex} ≃ 470 K for 1 ≤ J ≤ 5 by fitting the population densities of the rotational states to theoretical Boltzmann distributions. We conclude that UV pumping dominates the population of the higher rotational levels, as known from the H_{2} gas in the Milky Way.

∗ Data obtained under the DARA ASTRO-SPAS guest observer program; Research supported in part by the DARA

Accepted by: Astronomy & Astrophysics Letters

For preprints, contact deboer@astro.uni-bonn.de
Also available from http://xxx.lanl.gov/abs/astro-ph/9808256

ORFEUS II echelle spectra: H_{2} absorption in SMC gas

P. Richter^{1}, H. Widmann^{2}, K.S. de Boer^{1}, I. Appenzeller^{3}, J. Barnstedt^{2}, M. Gölz^{2}, M. Grewing^{4}, W. Gringel^{2}, N. Kappelmann^{2}, G. Krämer^{2}, H. Mandel^{3}, K. Werner^{2}

1 Sternwarte, Universität Bonn, D-53121 Bonn, Germany 2 Institut für Astronomie und Astrophysik, Abt. Astronomie, Universität Tübingen, D-72076 Tübingen, Germany 3 Landessternwarte Heidelberg, Königstuhl, D-69117 Heidelberg, Germany 4 IRAM, F-38406 Saint Martin d’Hères, France

We present a study of H_{2} in the SMC gas, based on space shuttle Far UV spectroscopy with ORFEUS and the Tü/HD echelle spectrograph, in the line of sight to the SMC star HD 5980. 17 absorption lines from the Lyman band have been analysed. Our line of sight crosses two clouds within the SMC.

We detect a cool molecular component near +120 km s^{−1}, where the H_{2} from the lowest 3 rotational states (J ≤ 2) is found. For this cloud we derive an excitation temperature of ≃ 70 K, probably the kinetic temperature of the gas. The cloud is located in the SMC foreground.

Another SMC component is visible at +160 km s^{−1}. Here we find unblended H_{2} absorption lines from levels 5 ≤ J ≤ 7. For this component we obtain an equivalent excitation temperature T_{ex} > 2350 K and conclude that this cloud must be highly excited by strong UV radiation from its energetic environment.

(Research supported in part by the DARA)

Accepted by: Astronomy & Astrophysics Letters

For preprints, contact prichter@astro.uni-bonn.de
Also available from http://xxx.lanl.gov/abs/astro-ph/9808258

Structure and Kinematics of the Interstellar Medium in front of SN1987A

J. Xu^{1} and A.P.S. Crotts

Columbia University, Dept. of Astronomy, 550 W. 120th St., New York, NY 10027, USA

High resolution (10 km/s) [N II] echelle spectra, sampled every 13 arcsec in a 6 arcmin by 6 arcmin region around SN1987A were obtained on the CTIO 4m telescope. The map shows a complicated velocity structure consistent with that reported previously for the interstellar medium. Three components, heliocentric velocity V_{hel} = 265, 277 and 285 km/s are identified as N157C. The radius of this large superbubble was found to expand at 10 km/s, with a lifetime of 6 million years and a total energy of 3 × 10^{51} ergs determined from its radius and velocity according to superbubble theory. The V_{hel} = 235 km/s component corresponds to the near side of 600 pc giant superbubble reported earlier. This bubble is over 10 million years old, and has blown out of the LMC disk. Two other components, V_{hel} = 255 and 245 km/s are identified as the inner major light echo ring (a double-shell structure) at about 130 pc in front of SN1987A. There are also two high velocity components, 300 and 313 km/s, which are possibly the far side of a superbubble in which SN1987A exploded. We also notice that there are two components at 269 and 301 km/s within 20 arcsec of SN1987A. These structures are probably due to the emission from the progenitor star’s red supergiant wind. We find that the time it took the SN1987A progenitor to move to the current location 300 pc behind N157C is comparable to the age of N157C as well as that of the progenitor itself.

Accepted by: The Astrophysical Journal

For preprints, contact arlin@astro.columbia.edu
Also available from the URL ftp://carmen.phys.columbia.edu/pub/arlin/astro-ph
or by anonymous ftp at carmen.phys.columbia.edu (IP# 128.59.196.23, in /pub/arlin/astro-ph

Some Characteristics of Current Star Formation in the 30 Doradus Nebula Revealed by HST/NICMOS

Nolan R. Walborn^{1}, Rodolfo H. Barbá^{1}, Wolfgang Brandner^{2}, Mónica Rubio^{3}, Eva K. Grebel^{4}, Ronald G. Probst^{5}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, Maryland 21218, USA 2 Jet Propulsion Laboratory/IPAC, California Inst. of Technology, 770 S. Wilson Ave, Pasadena, California 91125, USA 3 Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile 4 Lick Observatory, University of California, Santa Cruz, California 95064, USA 5 Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatories, Casilla 603, La Serena, Chile

The extensive “second generation” of star formation within the 30 Doradus Nebula, evidently triggered by the R136 central cluster around its periphery, has been imaged with HST/NICMOS. Many new IR sources, including multiple systems, clusters, and nebular structures, are found in these images. Six of the NICMOS fields are described here, in comparison with the WFPC2 images of the same fields. Knots 1-3 of Walborn & Blades (early O stars embedded in dense nebular knots) are all found to be compact multiple systems. Knot 1 is shown to reside at the top of a massive dust pillar oriented directly toward R136, whose summit has just been removed, exposing the newborn stellar system. Knots 1 and 3 are also near the brightest IR sources in the region, while parsec-scale jet structures are discovered in association with Knots 2 and 3. The Knot 2 structures consist of detached, nonstellar IR sources aligned on either side of the stellar system, which are interpreted as impact points of a highly collimated, possibly rotating bipolar jet on the surrounding dark clouds; the H_{2}O maser found by Whiteoak et al. is also in this field. These outflows from young massive stars in 30 Dor are the first extragalactic examples of the phenomenon. In the field of the pillars south of R136, recently discussed in comparison with the M16 pillars by Scowen et al., a new luminous stellar IR source has been discovered. These results establish the 30 Doradus Nebula as a prime region in which to investigate the formation and very early evolution of massive stars and multiple systems. The theme of triggered formation within the heads of extensive dust pillars oriented toward R136 is strong. In addition, these results provide further insights into the global structure and evolution of 30 Doradus, which are significant in view of its status as the best resolved extragalactic starburst.

Submitted to: The Astronomical Journal

For preprints, contact walborn@stsci.edu or rbarba@stsci.edu
Also available from the URL http://www.stsci.edu/~rbarba/publications.html#preprints
or by anonymous ftp at ftp.stsci.edu, /outside-access/out.going/rbarba

Flash ionization of the partially ionized wind of the progenitor of SN 1987A

P. Lundqvist

Stockholm Observatory, SE-133 36 Saltsjöbaden, Sweden

The H II region created by the progenitor of SN 1987A was further heated and ionized by the supernova flash. Prior to the flash, the temperature of the gas was ∼ 4000 − 5000 K, and helium was neutral, while the post-flash temperature was only slightly less than ∼ 10^{5} K, with the gas being ionized to helium-like ionization stages of C, N and O. We have followed the slow post-flash cooling and recombination of the gas, as well as its line emission, and find that the strongest lines should be N V λ1240 and O VI λ1034. Both these lines are good probes for the density of the gas, and suitable instruments to detect the lines are STIS on HST and FUSE, respectively. Other lines which may be detectable are N IV] λ1486 and [O III] λ5007, though they are expected to be substantially weaker. The relative strength of the oxygen lines is found to be a good tracer of the color temperature of the supernova flash. From previous observations, we put limits on the hydrogen density, n_{H}, of the H II region. The early N V λ1240 flux measured by IUE gives an upper limit which is n_{H} ∼ 180 η^{−0.40} cm^{−3}, where η is the filling factor of the gas. The recently reported emission in [O III] λ5007 at 2500 days requires n_{H} = (160 ± 12) η^{−0.19} cm^{−3}, for a supernova burst similar to that in the 500full1 model of Ensman & Burrows (1992). For the more energetic 500full2 burst the density is n_{H} = (215 ± 15) η^{−0.19} cm^{−3}. These values are much higher than in models of the X-ray emission from the supernova (n_{H} ∼ 75 cm^{−3}), and it seems plausible that the observed [O III] emission is produced primarily elsewhere than in the H II region. We also discuss the type of progenitor consistent with the H II region. In particular, it seems unlikely that its spectral type was much earlier than B2 Ia.

Accepted by: The Astrophysical Journal (Main Journal) For preprints, contact peter@astro.su.se

Also available from the URL ftp://www.astro.su.se/pub/supernova/preprints.html

WFPC2 observations of Star Clusters in the Magellanic Clouds. II. The Oldest Star Clusters in the Small Magellanic Cloud

K.J. Mighell^{1}, A. Sarajedini^{2}, and R.S. French^{3}

1 Kitt Peak National Observatory, National Optical Astronomy Obs., P. O. Box 26732, Tucson, AZ 85726-6732, USA 2 Dept. of Physics and Astronomy, San Francisco State Univ., 1600 Holloway Avenue, San Francisco, CA 94132, USA 3 Middle Tennessee State University, Physics & Astronomy Dept., WPS 219, P. O. Box 71, Murfreesboro, TN 37132, USA

We present our analysis of archival Hubble Space Telescope Wide Field Planetary Camera 2 (WFPC2) observations in F450W (∼B) and F555W (∼V) of the intermediate-age populous star clusters NGC 121, NGC 339, NGC 361, NGC 416, and Kron 3 in the Small Magellanic Cloud. We use published photometry of two other SMC populous star clusters, Lindsay 1 and Lindsay 113, to investigate the age sequence of these seven populous star clusters in order to improve our understanding of the formation chronology of the SMC. We analyzed the V vs B−V and M_{V} vs (B−V)_{o} color-magnitude diagrams of these populous Small Magellanic Cloud star clusters using a variety of techniques and determined their ages, metallicities, and reddenings. These new data enable us to improve the age-metallicity relation of star clusters in the Small Magellanic Cloud. In particular, we find that a closed-box continuous star-formation model does not reproduce the age-metallicity relation adequately. However, a theoretical model punctuated by bursts of star formation is in better agreement with the observational data presented herein.

Accepted by: The Astronomical Journal For preprints, contact mighell@noao.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9808091,
   or for full-resolution figures use: http://www.noao.edu/noao/staff/mighell/oldsmc/

Evolutionary synthesis of stellar populations: a modular tool

Claudia Maraston

Department of Astronomy, University of Bologna, Italy

A new tool for the Evolutionary Synthesis of Stellar Populations (EPS) is presented, which is based on three independent matrices, giving respectively: 1) the fuel consumption during each evolutionary phase as a function of stellar mass; 2) the typical temperatures and gravities during such phases; 3) colours and bolometric corrections as functions of gravity and temperature. The modular structure of the code allows to easily assess the impact on the synthetic spectral energy distribution of the various assumptions and model ingredients, such as, for example, uncertainties in stellar evolutionary models, mixing length, the temperature distribution of horizontal branch (HB) stars, AGB mass loss, and colour-temperature transformations. The so-called “AGB-Phase Transition” in Magellanic Cloud clusters is used to calibrate the contribution of the Thermally Pulsing Asymptotic Giant Branch phase (TP-AGB) to the synthetic integrated luminosity. As an illustrative example, solar metallicity (Y = 0.27, Z = 0.02) models, with ages ranging between 30 Myr and 15 Gyr and various choices for the slope of the Initial Mass Function (IMF), are presented. Synthetic broad band colours and the luminosity contributions of the various evolutionary stages are compared with LMC and Galactic globular cluster data. In all these cases, a good agreement is found. Finally, we show the evolution of stellar mass-to-light ratios in the bolometric and U,B,V,R, and K passbands, in which the contribution of stellar remnants is accounted for.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact maraston@usm.uni-muenchen.de
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9807338

The Optical Gravitational Lensing Experiment.

Photometry of the MACHO-SMC-98-1 Binary Microlensing Event

A. Udalski^{1}, M. Kubiak^{1}, M. Szymański^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present photometry of the unique binary microlensing event MACHO-SMC-98-1 collected by the OGLE group. Particularly interesting observation was made close to the first caustic crossing which was not covered by observations of other groups. It allows to test proposed models of which Model 1 proposed by PLANET group seems to be in the best agreement with the OGLE observations.

Accepted by: Acta Astronomica 1998, 48

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9808077

Optimal Microlensing Observations

Andrew P. Gould

Dept. of Astronomy, Ohio State University, Columbus, OH 43210-1106, USA

One of the major limitations of microlensing observations toward the Large Magellanic Cloud (LMC) is the low rate of event detection. What can be done to improve this rate? Is it better to invest telescope time in more frequent observations of the inner high surface-brightness fields, or in covering new, less populated outer fields? How would a factor 2 improvement in CCD sensitivity affect the detection efficiency? Would a series of major (factor 2–4) upgrades in telescope aperture, seeing, sky brightness, camera size, and detector efficiency increase the event rate by a huge factor, or only marginally? I develop a simplified framework to address these questions. With observational resources fixed at the level of the MACHO and EROS experiments, the biggest improvement (factor ∼2) would come by reducing the time spent on the inner ∼ 25 deg^{2} and applying it to the outer ∼ 100 deg^{2}. By combining this change with the characteristics of a good medium-size telescope (2.5 m mirror, 1′′ point spread function, thinned CCD chips, 1 deg^{2} camera, and dark sky), it should be possible to increase the detection of LMC events to more than 100 per year (assuming current estimates of the optical depth apply to the entire LMC).

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9807350

Abstracts of Non-Refereed Papers

The Formation and Evolution of LMC Globular Clusters: the Database

S.F. Beaulieu^{1}, R. Elson^{1}, G. Gilmore^{1}, R.A. Johnson^{1}, N. Tanvir^{1} and B. Santiago^{2}

1 Institute of Astronomy, University of Cambridge, Madingley Road CB3 0HA, England 2 Instituto de Fisica, Universidade Federal do Rio Grande do Sul, Av. Bento Goncalves 9500, Porto Alegre RS, CEP: 91501-970, Brazil

We present details of the database from a large Cycle 7 HST project to study the formation and evolution of rich star clusters in the Large Magellanic Cloud (see Elson et al., this volume). Our data set, which includes NICMOS, WFPC2 and STIS images of the cores and outer regions of 8 clusters, will enable us to derive deep luminosity functions for the clusters and to investigate the universality of the stellar initial mass function. We will look for age spreads in the youngest clusters, quantify the population of binary stars in the cores of the clusters and at the half-mass radii, and follow the development of mass segregation.

To appear in: IAU Symp. 190, New Views of the Magellanic Clouds, Eds. Chu et al., ASP Conference Series

For preprints, contact beaulieu@ast.cam.ac.uk
Also available from the URL http://www.ast.cam.ac.uk/LMC

Deep STIS Luminosity Functions for LMC Clusters

R. Elson, N. Tanvir, G. Gilmore, R.A. Johnson, & S. Beaulieu

Institute of Astronomy, Madingley Rd., Cambridge CB3 0HA England

We present deep luminosity functions derived from HST STIS data for three rich LMC clusters (NGC 1805, NGC 1868, and NGC 2209), and for one Galactic globular cluster (NGC 6553). All of the LMC cluster luminosity functions are roughly consistent with a Salpeter IMF or with the solar neighbourhood IMF from Kroupa, Tout & Gilmore (1993). They continue to rise at least to 0.7 M_{⊙}. NGC 1868 shows evidence for mass segregation which may be primordial. A comparison of deep luminosity functions for seven Galactic globulars shows that the luminosity functions are eroded at low masses by amounts that are strongly correlated with distance from the Galactic plane.

To appear in: IAU Symp. 190, New Views of the Magellanic Clouds, Eds. Chu et al., ASP Conference Series For preprints, contact elson@ast.cam.ac.uk

Also available from the URL http://www.ast.cam.ac.uk/LMC/

A 16 Millisecond X-Ray Pulsar in the Crab-Like SNR N157B: Fast Times at 30 Doradus

E.V. Gotthelf^{1}, W. Zhang^{1}, F.E. Marshall^{1}, J. Middleditch^{2}, Q.D. Wang^{3}

1NASA/Goddard Space Flight Center, Code 660.2, Greenbelt, MD 20771, USA 2 Los Alamos National Laboratory, MS B256, CIC-19, Los Alamos, NM 87545, USA 3 Dearborn Observatory, Northwestern University, 2131 Sheridan Road, Evanston, IL 60208, USA

The supernova remnant N157B (30 Dor B, SNR 0539-69.1, NGC 2060), located in the Tarantula Nebula of the Large Magellanic Cloud, has long been considered a possible Crab-like remnant. This hypothesis has been confirmed, quite spectacularly, with the discovery of PSR J0537−6910, the remarkable 16 ms X-ray pulsar in N157B. PSR J0537−6910 is the most rapidly spinning pulsar found to be associated with a supernova remnant. Here we report our discovery and summarize the properties of this pulsar and its supernova remnant.

To appear in: Special issue of Memorie della Societa’ Astronomica Italiana (Proceedings)

For preprints, contact gotthelf@gsfc.nasa.gov
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9808240

The Contribution of Microlensing Surveys to the Distance Scale

J.P.Beaulieu^{1} & W.J. de Wit^{2}

1 Kapteyn Instituut, Postbus 800, NL-9700 AV Groningen, The Netherlands 2 Universiteit Utrecht, Sterrekundig Instituut, Postbus 80000, NL-3508 TA Utrecht, The Netherlands

In the early nineties several teams started large scale systematic surveys of the Magellanic Clouds and the Galactic Bulge to search for microlensing effects. As a by product, these groups have created enormous time-series databases of photometric measurements of stars with a temporal sampling duration and accuracy which are unprecedented. They provide the opportunity to test the accuracy of primary distance indicators, such as Cepheids, RR Lyrae stars, the detached eclipsing binaries, or the luminosity of the red clump.

We will review the contribution of the microlensing surveys to the understanding of the physics of the primary distance indicators, recent differential studies and direct distance determinations to the Magellanic Clouds and the Galactic Bulge.

To appear in: ‘Post-Hipparcos Cosmic Candles’, A. Heck & F. Caputo (Eds), Kluwer Academic Publ., Dordrecht, in press.

For preprints, contact Beaulieu@astro.rug.nl, w.j.m.dewit@astro.uu.nl
Also available from the URL http://www.astro.rug.nl/~beaulieu/,
    and http://xxx.lanl.gov/abs/astro-ph/9808080